OpenStudy (anonymous):

15 divided by 3 + 3/8 *8/9

mathslover (mathslover):

\[\large{\frac{15}{3}+\frac{3}{8}\times \frac{8}{9}}\]

mathslover (mathslover):

is this the question @shivraj ?

OpenStudy (anonymous):

yes

mathslover (mathslover):

ok so here we go: \[\large{5+ \frac{3}{9}\times \frac{8}{8}}\] \[\large{5+\frac{1}{3}}\]

mathslover (mathslover):

that is 16/3 got it?